what is the value of the variable x and y in:
x-9y=-13
y=-x+7

Answer by mananth(16946)

x-9 y = -13 .............1

x+y = 7 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 9
1 x -9 y = -13
9 x + 9 y = 63
Add the two equations
10 x = 50
/ 10
x = 5
plug value of x in (1)
1 x -9 y = -13
5 -9 y = -13
-9 y = -13 -5
-9 y = -18
y = 2